one leg of a right isosceles triangle is 6ft. what is the area of the triangle? and how do you work it out

Answer by pmesler(52)

To figure out this problem you need to know two things: the definition of an isosceles triangle, and the formula to find the area of a triangle. Okay, so what's an isosceles triangle? An isosceles triangle is a triangle where two of its sides are equal. The area of a triangle is
A = 1/2 b * h. Where b and h are the lengths of the base and height, respectively. So, what's the height?
The height of any triangle is always the leg that is perpendicular to the base. In the problem it says that the triangle is a right isosceles triangle. Right away that tells us two things: Two sides of this triangle are going to be the same AND one of the legs must be perpendicular to another leg in order to form a 90 degree or right angle, hence the name right triangle.
Since you generally need to know the length of the height to find the area of a triangle, let's assume that the height is also 6ft. That means the base is 6ft and the height is 6ft. Finding the area will be a cinch. Just plug in the values for the base and height like so:
A = (6) * (6)/2
A = 36/2
A = 18ft^2. The area of the triangle is 18ft^2.